A spy camera is said to be able to read the numbers on a car's license plate. If the numbers on the plate are 6.80 cm apart, and the spy satellite is at an altitude of 180 km, what must be the diameter of the camera's aperture? (Assume light with a wavelength of 570 nm.)
Answer #1

from Raleigh criterion

= 1.22 L / D

also

= dx / dr

so considering both equation above

dx / dr = 1.22 L/ D

6.8* 10^-2 / 180000 = 1.22* 570* 10^-9 / D

D = 1.841 m

========